OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12

OpenStudy (anonymous):

If you divide by -3, you'll have |2x+6| = 4. We can also factor out a 2, and divide that on both sides, to have |x+3|=2. Now it's really easy. The absolute value bars mean that x+3 could either be a positive 2 or a -2 and it would still satisfy the equation. if x+3 = -2, then x = -5. if x+3 = 2, then x = -1.

OpenStudy (anonymous):

thank you